30. Suppose that in an economy, output per worker and capital per worker grew by 2.42%, and 2.48%, respectively. If the share of capital in output is 0.25:

(a) Calculate the total factor productivity.

(b) Is there a reason to believe that this economy is at steady state?


1
Expert's answer
2022-01-11T09:30:54-0500

a) "A=\frac{Y}{K^\alpha L^\beta}"

Where;

Y= Total output

A- Total factor productivity

K= Capital input

L=labor Input

"\alpha , \beta =Contribution"


"A=\frac{1.0242}{1.0248^\frac{1}{4} 1^\frac{3}{4}}"

= 1.01795

= 101.8%

b) There is a reason to say that the economy is stable because its total factor productivity is more than 100% hence showing how stable the economy is.